OpenStudy (anonymous):

A falcon flew 70 miles in 3 hours. What was the speed of the falcon? speed = ; 1 mile = 1609 meters. (2 points) 10.4 meters per second 23.3 meters per second 52.2 meters per second 93.9 meters per second

OpenStudy (anonymous):

speed = distance/time s = 70/3 s = 23.33 miles per hour 1 mile = 1609 meters so 23.33 miles = (1609 x 23.33) 37537.97 meters 1 hour = 60 minutes = 60 sec per min = (60 x 60) = 3600 seconds 37537.97/3600 = 10.4 meters per second :)